Inorganic Chemistry

What is ∆S of a cell if Ecell=0.045V at 298K ang temperature coefficient of cell is 3×10-4VK-1

Karan Kumawat
7 Years agoGrade

1 Answer

Khimraj
7 Years ago
According to Gibbs helmholtz equation , 
\bold{\Delta{G}=\Delta{H}+T(\frac{\delta(\Delta{G})}{\delta{T}})_P}
Here ∆G is Gibbs free energy , ∆H is heat ,T is temperature and  is rate of change of Gibbs free energy at constant pressure.
But we know, ∆G = -nEF 
∴  \bold{(\frac{\delta(\Delta{G})}{\delta{T}})_P} = -nF\bold{(\frac{\delta{E}}{\delta{T}})_P}
Put it in above equation, 
Then, ∆G = ∆H - nFT\bold{(\frac{\delta{E}}{\delta{T}})_P}
Now, compare this equation with , ∆G = ∆H - T∆S, 
Here ∆S is the entropy change.
Hence, ∆S = - nF\bold{(\frac{\delta{E}}{\delta{T}})_P}
,here is temperature coefficient and ∆S is entropy change.